Suppose a recurrence relation




an=4an−1−4an−2



where a1=14 and a2=40




can be represented in explicit formula, either as:




Formula 1:



an=pxn+qnxn



or




Formula 2:



an=pxn+qyn






where



x



and



y



are roots of the characteristic equation.







Determine



p



and



q



.




1
Expert's answer
2021-11-26T13:43:46-0500
"a_n-4a_{n\u22121}+4a_{n\u22122}=0"

characteristic equation:


"r^2-4r+4=0"

"(r-2)^2=0"

"r=2" is the only characteristic root. Therefore we know that the solution to the recurrence relation has the form


"a_n=a2^n+bn2^n"

Given "a_1=14 ,a_2=40"


"a_1=a(2^1)+b(1)(2^1)=14"

"a_2=a(2^2)+b(2)(2^2)=40"

"a+b=7""\\

 a+2b=10"


"a=4""b=3"

Therefore the solution to the recurrence relation is


"a_n=4\cdot2^n+3n\cdot2^n"